Find all real or imaginary solutions to the equation.

49x^2+9 = 42x

Answer by rapaljer(4671)

49x^2+9 = 42x

First set the equation equal to zero and place in descending powers of x:
49x^2 - 42x + 9 = 0

Factor it. (NOTE: It just happens to factor!!)
{7x - 3)(7x - 3) = 0
7x = 3
x= 3/7

R^2 at SCC